# Everything Everywhere All At Once: A First Principles Numerical Demonstration of Emergent Decoherent Histories

### v2

@@ Abstract
Within the histories formalism the decoherence functional is a formal tool to investigate the emergence of classicality in isolated quantum systems, yet an explicit evaluation of it from first principles has not been reported. We provide such an evaluation for up to five-time histories based on exact numerical diagonalization of the Schrödinger equation. We find a robust emergence of decoherence for slow and coarse observables of a generic random matrix model and extract a finite size scaling law by varying the Hilbert space dimension over four orders of magnitude. Specifically, we conjecture and observe an exponential suppression of coherent effects as a function of the particle number of the system. This suggests a solution to the preferred basis problem of the many worlds interpretation (or the set selection problem of the histories formalism) within a minimal theoretical framework—without relying on environmentally induced decoherence, quantum Darwinism, Markov approximations or ensemble averages. We further discuss the implications of our results for the wave function of the Universe, interpretations of quantum mechanics and the arrow(s) of time.
@@ I Introduction
Here, we use the decoherence functional (DF) introduced within the consistent or decoherent histories framework (or simply “histories framework” for short) Griffiths (1984); Gell-Mann and Hartle (1990); Omnès (1992); Dowker and Halliwell (1992); Gell-Mann and Hartle (1993); Halliwell (1995); Dowker and Kent (1996); Griffiths (2002); Gell-Mann and Hartle (2007); Griffiths (2019) to rigorously investigate the emergence of multi-time classicality (see Sec. II for the technical details). We numerically show that bases defined by a slow and coarse observable of a non-integrable system are robustly decoherent according to this criterion. More specifically, we show for up to five-time histories and a Hilbert space dimension $D$ varying over four orders of magnitude that quantum effects (suitably quantified below) are exponentially suppressed as a function of the particle number $N$ of the system (with $N\sim\log D$). This provides a firm starting point to discuss the MWI and to address the set selection problem of the histories framework Paz and Zurek (1993); Dowker and Kent (1995, 1996); Riedel et al. (2016). We further discuss important cosmological implications, namely that the emergence of classicality is ubiquituous, that almost all initial wave functions can give rise to interesting universes, and that the branching of the wave function is a priori not related to any arrow of time.
@@
Perhaps surprisingly, our results do not rely on environmentally induced decoherence (EID) Zurek (2003); Joos et al. (2003); Schlosshauer (2019) [^1] Note the unfortunate double meaning of decoherence in the literature:a priori there is no connection between the DF in the histories framework and the concept of EID. or its refinement to quantum Darwinism Zurek (2009); Korbicz (2021); Zurek (2022). This is surprising because the widely proclaimed (sole) answer to the question why the MWI gives rise to classically looking universes is EID, and many researchers worked on connecting the MWI or the histories framework to EID and quantum Darwinism Finkelstein (1993); Saunders (1993); Paz and Zurek (1993); Diósi et al. (1995); Brun (1997); Yu (1998); Vaidman (1998); Brun (2000); Riedel et al. (2016); Albrecht et al. (2022); Touil et al. (2022); Strasberg (2023), yet an explicit evaluation of the DF following the desiderata (A) and (B) is still missing. Only if one invokes a rigorous notion of multi-time quantum Markovianity Pollock et al. (2018); Li et al. (2018); Milz and Modi (2021), clear connections between the DF and EID have been established Paz and Zurek (1993); Diósi et al. (1995); Brun (1997); Yu (1998); Brun (2000); Strasberg (2023), but this only shifts the problem of proving multi-time decoherence to proving multi-time Markovianity, which is a daunting task too Dümcke (1983); Ford and O’Connell (1996); Figueroa-Romero et al. (2019, 2021); Strasberg et al. (2023).
@@
In contrast, the present approach does not rely on any system-environment tensor product splitting, although it is important to emphasize that it is not in conflict with EID or quantum Darwinism when applied to it. Instead, we only consider slow and coarse observables of isolated, non-integrable quantum systems, similar to the approach taken by van Kampen in 1954 Van Kampen (1954). While the importance of slow and coarse (or quasi-conserved) observables for the emergence of decoherence has been anticipated in the histories formalism Gell-Mann and Hartle (1993); Brun and Halliwell (1996); Halliwell (1998, 1999a); Calzetta and Hu (1999); Halliwell (2003); Gell-Mann and Hartle (2007); Halliwell (2010), non-integrability has never been considered a key factor by proponents of the histories, EID or quantum Darwinism framework: van Kampen’s work remained ignored despite him emphasizing its importance for the quantum-to-classical transition van Kampen et al. (2000); van Kampen (2008). Only recently rigorous evidence for van Kampen’s ideas slowly accumulated Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Strasberg et al. (2023); Strasberg (2023). Still, it has not yet been considered in light of the MWI and a first principles evaluation of the DF remains missing.
@@
To summarize, our objective is to subject the MWI and the question whether it is compatible with our perceived classical reality to a rigorous, quantitative test based on the minimal desiderata (A) and (B). To this end, we introduce the general theoretical framework to address this question in Sec. II. Then, we present extensive numerical results for a heat exchange model (an archetypical example of a nonequilibrium process in thermodynamics) in Sec. III. Section IV discusses important implications and we conclude and provide perpectives in Sec. V.
@@ II.2 Physical meaning and further terminology
We continue by commenting on the connection between the DHC and the MWI as we view it in this work. As emphasized in the introduction, the definition of a “world” within the MWI is a priori complicated by the existence of a continuum of mathematically conceivable different worlds (preferred basis problem). Moreover, even among the proponents of the MWI there is no universally agreed on quantitative definition of a world. Here, we use this liberty and interpret the DHC as a minimal criterion to define a world within the MWI. This is motivated by what we said above: the classical world that we perceive is compatible with quantum mechanics precisely when the observables we look at satisfy the DHC. We believe this view is in unison with proponents of the MWI, for instance, Vaidman verbally defines a world as “the totality of macroscopic objects […] in a definite classically described state” Vaidman (2021). Moreover, a close connection between the DHC and the existence of records has been established by various researchers Albrecht (1992); Gell-Mann and Hartle (1993); Finkelstein (1993); Paz and Zurek (1993); Halliwell (1999b); Dodd and Halliwell (2003); Riedel et al. (2016); Hartle (2016).
@@ II.4 Physical origin of decoherence
The physical origin of decoherence has been discussed at many places and shall not be repeated here in detail. We therefore limit ourselves to listing the four assumptions used by van Kampen Van Kampen (1954) (see also Refs. Strasberg et al. (2023); Strasberg (2023) for a more recent discussion) followed by a perhaps oversimplified yet hopefully intuitive analogy.
The four assumptions are the following. First, one needs a large Hilbert space dimension $D$, which is certainly satisfied for macroscopic systems. Typically, $D={\cal{O}}(10^{N})$ and $N={\cal{O}}(10^{23})$ to give some numbers. Second, one must consider a coarse coarse-graining satisfying $M\ll D$. As discussed above, this is satisfied for all that we humans can perceive. Third, the considered projectors must evolve slowly in a suitable sense such that the unobserved microscopic degrees of freedom have time to self-average or randomize. Common examples include hydrodynamic modes, collective degrees of freedom, weakly coupled subsystems, or more generally projectors that almost commute with the total Hamiltonian. Fourth, the isolated system should be non-integrable.
This last point is certainly still subject to debate, not the least because a universally agreed on definition of quantum (non-)integrability is lacking. For instance, van Kampen assumed that the energy gap spectrum of the Hamiltonian $H$ is non-degenerate (apart from rare accidental degeneracies) Van Kampen (1954), the results of Ref. Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Strasberg et al. (2023) were based on the eigenstate thermalization hypothesis Srednicki (1999); D’Alessio et al. (2016); Deutsch (2018), and here we use random matrix theory as also done in Ref. Nation and Porras (2020); Strasberg (2023). Since already classical systems tend to be chaotic (e.g., the Newtonian three-body problem), we believe that non-integrability can hardly count as an assumption in our Universe, even though it is an open question whether it is strictly necessary.
@@ III.1 Setup
Each of these blocks is in principle fully determined by the microscopic Hamiltonian (12), but we only want to capture four main features here. First, the systems $A$ and $B$ are assumed to be identical from a thermodynamic point of view. In particular, we assume that the relation between energy and temperature is the same in $A$ and $B$ such that their energy difference is proportional to their temperature difference. Second, we consider normal thermodynamic systems where the subspace corresponding to an equal energy distribution (or the same temperature according to the previous agreement) is the largest “equilibrium” subspace, i.e., $V_{0}>V_{-},V_{+}$. Third, we assume a weak interaction between $A$ and $B$, i.e., the interaction energy $\lambda H_{I}$ is supposed to be negligible (otherwise it would not be justified to restrict the discussion to a microcanonical energy window defined by the sum $E_{A}+E_{B}$ of local energies only). Fourth, we want to mimic the interaction of two generic complex (and thus non-integrable) many-body systems. According to common knowledge in statistical mechanics, this can be efficiently done by using random matrix theory Wigner (1967); Brody et al. (1981); Deutsch (1991); Beenakker (1997); Guhr et al. (1998); Haake (2010); D’Alessio et al. (2016); Deutsch (2018); Reimann and Dabelow (2021).
@@ III.2 Results
To continue with the many worlds simulation, we now choose an equilibrium initial state with $p_{x}(0)=V_{x}/D$, for reasons that will also become clear in Sec. IV. Naively, one would expect nothing interesting in such a Universe as indicated in Fig. 3(b): aside from fluctuations, which are exponentially suppressed as a function of the Hilbert space dimension, nothing seems to happen. But interestingly, a completely different picture is possible: even from an equilibrium state many different interesting nonequilibrium universes can emerge. However, to warrant such a conclusion, we first of all need to ensure that the histories defined by the present coarse-graining are decoherent, i.e., they obey the DHC condition as discussed in the previous section. In the following, the evaluation of the DF is done for constant time intervals $t_{k+1}-t_{k}$ equal to the nonequilibrium relaxation time scale $\tau$ defined in Eq. (17).
@@
The picture that emerges from our model is depicted in Fig. 6. Figure 6 (a) shows a branching tree structure with respect to histories defined by energetic macrostates labeled by $x\in\{-,0,+\}$ for $L=3$ time steps. Importantly, we can now reason about these histories using a classical ontological model for sufficiently large $D$. To facilitate talking about it, we use Boltzmann’s entropy concept $S_{B}(x)=k_{B}\ln V_{x}$. We label transitions of the system from a lower to a higher Boltzmann entropy state (i.e., from $+$ or -- to 00) as “forward” transitions (blue dashed lines) to indicate that they comply with the conventional arrow of time when $t_{0}$ denotes the initial time. However, “backward” transitions (pink dash-dotted lines) from a higher to a lower entropy state (i.e., from 00 to $+$ or --) are also possible. Finally, transitions involving no change in Boltzmann entropy are labeled by “no arrow” (black lines). The Multiverse thus consists of different histories or branches that do not interfere and that describe universes with different arrows of time (including the possibility of universes with no arrow of time and both arrows of time) determined by the question whether heat flows from hot to cold or vice versa. For one realization of the Hamiltonian and one realization of the initial state we depict the probability for each history, which can be now interpreted classically, by the histogram below the tree-like branching structure in Fig. 6 (b). As a consistency check, we immediately see that the history $\textbf{x}=(0,0,0)$, in which the universe resides in the dominant equilibrium macrostate $x=0$, is the most probable one. Note that the probability for this history is different from the probability $(V_{0}/D)^{3}=0.6^{3}=0.216$ that one would expect by sampling the histories at time intervals comparable (or larger) to the equilibration time. The time steps $\tau$ we use are therefore clearly in the nonequilibrium regime.
@@
An important consistency check is to ask about the probability of forward or backward arrows of time. Since the Schrödinger equation obeys time reversal symmetry and since the initial equilibrium state that we chose also does not introduce any time asymmetry, we should expect to get a time symmetric answer in this case—a point to which we return from a broader perspective in Sec. IV.4. The answer is shown in Fig. 6 (c), (d) and (e) where we plot the probability for a forward arrow of time (blue), no arrow of time (black) and a backward arrow of time (pink) for histories of length $L=3$, $L=4$ and $L=5$, respectively. Here, histories with multiple arrows of time contribute according to their “net” arrow of time, i.e., the history $\textbf{x}=(0,+,0)$ has no overall arrow of time. Owing to the fact that we have only three macrostates, there can be no histories with two or more net forward or backward arrows of time. To conclude, we see that the arrows are indeed symmetrically distributed in our model as it should be. Some non-visible small random fluctuations (in the third significant digit) are due to the fact that a random initial state can still have a small preference to evolve into a higher or lower entropy region, i.e., it is not perfectly symmetric under time reversal. The histograms in Fig. 6 are based on a single realization of the Hamiltonian and initial state, but we again found this behaviour to be typical.
@@
The results for the emergence of decoherence with initial energy eigenstates are shown in Figs. 7 and 8. We again observe a rather typical behaviour and a scaling law suggesting the exponential suppression of quantum effects as a function of the particle number. However, compared with Figs. 4 and 5 the variance in the datapoints is appreciably larger, i.e., there is less typicality. Moreover, the exponents are different. For $\epsilon$ we find $\alpha\approx 0.44$, which is close to the previous value of $\alpha\approx 0.5$. Instead, for $\Delta_{L}^{\text{max}}$ we find $\alpha\approx 0.2$, which is half of the value we found for Haar random initial equilibrium states. The preliminary conclusions that we can draw from these data is that correlation functions of energy eigenstates behave differently from Haar random equilibrium states (which will typically overlap with many energy eigenstates), which is an interesting result in its own right. In particular, Albrecht et al. recently introduced an “eigenstate einselection hypothesis” (Appendix B in Ref. Albrecht et al. (2022)), where they claim that the “general features of the consistent histories quantities are unchanged” for energy eigenstates. On a quantitative level we see here clear evidence for the opposite, but from a qualitative point of view we observe the same: the difference between $D^{-0.4}$ and $D^{-0.2}$ might not matter in practice, say, for $D\geq 10^{100}$. Thus, the emergence of decoherence as studied here seems to be a rather robust phenomenon if we restrict the attention to slow and coarse observables of a non-integrable quantum many-body system.
@@
The result is a very different behaviour of decoherence as seen in Figs. 9 and 10. While typicality still holds well, the previously observed exponential suppression of quantum effects ceases to exist for histories of length $L\geq 3$. Instead, the scaling exponent of $\epsilon$ is around $\alpha=0.15$ in Fig. 9 and even negative values are possible for $\Delta_{L}^{\text{max}}$ as shown in Fig. 10 for $L=5$. Recalling that we have to base our conclusions on a finite (and rather small) amount of samples, a clear-cut conclusion seems not possible: exponential suppression of quantum effects, which requires $\alpha>0$, seems possible but is not warranted, and it is certainly much weaker than in the weak coupling (slow observable) regime.
@@
Notably, exponential suppression of quantum effects (with a large exponent $\alpha\approx 0.5$) still holds for the shortest length of histories with $L=2$. This seems counterintuitive, but it can be explained by recalling that we have chosen a Haar random initial equilibrium state of the form (18) with $p_{x}(0)=V_{x}/D$ in Figs. 9 and 10. The two-time DF then only probes quantum features of a two-time equilibrium correlation function. They are decoherent because, first, $\mathbb{E}[\Pi_{x_{0}}|\psi(t_{0}){\rangle\langle}\psi(t_{0})|\Pi_{y_{0}}]\sim\delta_{x_{0},y_{0}}$ where $\mathbb{E}[\dots]$ denotes a Haar random average, and second, typicality implies that most pure states $|\psi(t_{0})\rangle$ behave similarly. Indeed, for this situation an analytical proof of decoherence has been given in Ref. Strasberg et al. (2023).

### v4

# Everything Everywhere All At Once: First Principles Numerical Demonstration of Emergent Decoherent Histories

@@ Abstract
Within the histories formalism the decoherence functional is a formal tool to investigate the emergence of classicality in isolated quantum systems, yet an explicit evaluation of it from first principles has not been reported. We provide such an evaluation for up to five-time histories based on exact numerical diagonalization of the Schrödinger equation. We find a robust emergence of decoherence for slow and coarse observables of a generic random matrix model and extract a finite size scaling law by varying the Hilbert space dimension over four orders of magnitude. Specifically, we conjecture and observe an exponential suppression of coherent effects as a function of the particle number of the system. This suggests a solution to the preferred basis problem of the many worlds interpretation (or the set selection problem of the histories formalism) within a minimal theoretical framework—without relying on environmentally induced decoherence, quantum Darwinism, Markov approximations, low-entropy initial states or ensemble averages.
@@ I Introduction
Here, we use the decoherence functional (DF) introduced within the consistent or decoherent histories framework (or simply “histories framework” for short) Griffiths (1984); Gell-Mann and Hartle (1990); Omnès (1992); Dowker and Halliwell (1992); Gell-Mann and Hartle (1993); Halliwell (1995); Dowker and Kent (1996); Griffiths (2002); Gell-Mann and Hartle (2007); Griffiths (2019) to rigorously investigate the emergence of multi-time classicality (see Sec. II for the technical details). We numerically show that bases defined by a slow and coarse observable of a non-integrable system are robustly decoherent according to this criterion. More specifically, we show for up to five-time histories and a Hilbert space dimension $D$ varying over four orders of magnitude that quantum effects (suitably quantified below) are exponentially suppressed as a function of the particle number $N$ of the system (with $N\sim\log D$). This provides a firm starting point to discuss the MWI and to address the set selection problem of the histories framework Paz and Zurek (1993); Dowker and Kent (1995, 1996); Riedel et al. (2016). Our results also explicitly show that the emergence of classicality is ubiquituous, that almost all initial wave functions can give rise to interesting universes, and that the branching of the wave function is a priori not related to any arrow of time.
@@
Perhaps surprisingly, our results do not rely on environmentally induced decoherence (EID) Zurek (2003); Joos et al. (2003); Schlosshauer (2019) [^1] Note the unfortunate double meaning of decoherence in the literature:a priori there is no connection between the DF in the histories framework and the concept of EID. or its refinement to quantum Darwinism Zurek (2009); Korbicz (2021); Zurek (2022). This is surprising because the widely proclaimed (sole) answer to the question why the MWI gives rise to classically looking universes is EID, and many researchers worked on connecting the MWI or the histories framework to EID and quantum Darwinism Finkelstein (1993); Saunders (1993); Paz and Zurek (1993); Diósi et al. (1995); Brun (1997); Yu (1998); Vaidman (1998); Brun (2000); Riedel et al. (2016); Riedel (2017); Albrecht et al. (2022); Touil et al. (2022); Strasberg (2023), yet an explicit evaluation of the DF following the desiderata (A) and (B) is still missing. Only if one invokes a rigorous notion of multi-time quantum Markovianity Pollock et al. (2018); Li et al. (2018); Milz and Modi (2021), clear connections between the DF and EID have been established Paz and Zurek (1993); Diósi et al. (1995); Brun (1997); Yu (1998); Brun (2000); Strasberg (2023), but this only shifts the problem of proving multi-time decoherence to proving multi-time Markovianity, which is a daunting task too Dümcke (1983); Ford and O’Connell (1996); Figueroa-Romero et al. (2019, 2021); Strasberg et al. (2023).
@@
In contrast, the present approach does not rely on any system-environment tensor product splitting, although it is important to emphasize that it is not in conflict with EID or quantum Darwinism when applied to it. Instead, we only consider slow and coarse observables of isolated, non-integrable quantum systems, similar to the approach taken by van Kampen in 1954 Van Kampen (1954). While the importance of slow and coarse (or quasi-conserved) observables for the emergence of decoherence has been anticipated in the histories formalism Gell-Mann and Hartle (1993); Brun and Halliwell (1996); Halliwell (1998, 1999a); Calzetta and Hu (1999); Halliwell (2003); Gell-Mann and Hartle (2007); Halliwell (2010), non-integrability has never been considered a key factor by proponents of the histories, EID or quantum Darwinism framework: van Kampen’s work remained ignored despite him emphasizing its importance for the quantum-to-classical transition van Kampen et al. (2000); van Kampen (2008). More precisely, non-integrability is here taken to mean that random matrix theory captures well the relevant dynamic behaviour. While this is known to be true in many situations Wigner (1967); Brody et al. (1981); Beenakker (1997); Guhr et al. (1998); Haake (2010); D’Alessio et al. (2016); Deutsch (2018), in the context of the quantum-to-classical transition rigorous evidence for van Kampen’s idea only slowly accumulated recently Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Strasberg et al. (2023); Strasberg (2023). Moreover, it has not yet been considered in light of the MWI and a first principles evaluation of the DF remains missing.
@@
To summarize, our objective is to subject the MWI and the question whether it is compatible with our perceived classical reality to a rigorous, quantitative test based on the minimal desiderata (A) and (B). To this end, we introduce the general theoretical framework to address this question in Sec. II. Then, we present extensive numerical results for a heat exchange model (an archetypical example of a nonequilibrium process in thermodynamics) in Sec. III. Section IV discusses the set selection problem in view of our findings, and we conclude and provide perpectives in Sec. V.
@@ II.2 Physical meaning and further terminology
We continue by commenting on the connection between the DHC and the MWI as we view it in this work. As emphasized in the introduction, the definition of a “world” within the MWI is a priori complicated by the existence of a continuum of mathematically conceivable different worlds (preferred basis problem). Moreover, even among the proponents of the MWI there is no universally agreed on quantitative definition of a world (see, e.g., Ref. Riedel (2023) for a short overview). Here, we use this liberty and interpret the DHC as a minimal criterion to define a world within the MWI. This is motivated by what we said above: the classical world that we perceive is compatible with quantum mechanics precisely when the observables we look at satisfy the DHC. We believe this view is in unison with proponents of the MWI, for instance, Vaidman verbally defines a world as “the totality of macroscopic objects […] in a definite classically described state” Vaidman (2021). Moreover, only past “events” giving rise to decoherent histories can leave records about having “happened” in the present state of the Universe, which highlights the importance of the DHC Albrecht (1992); Gell-Mann and Hartle (1993); Finkelstein (1993); Paz and Zurek (1993); Halliwell (1999b); Dodd and Halliwell (2003); Riedel et al. (2016); Hartle (2016).
@@ II.4 Physical origin of decoherence
The physical origin of decoherence has been discussed at many places and shall not be repeated here in detail. We therefore limit ourselves to listing the four assumptions used by van Kampen Van Kampen (1954) followed by a perhaps oversimplified yet hopefully intuitive analogy.
The four assumptions are the following. First, one needs a large Hilbert space dimension $D$, which is certainly satisfied for macroscopic systems. Typically, $D={\cal{O}}(10^{N})$ and $N={\cal{O}}(10^{23})$ to give some numbers. Second, one must consider a coarse coarse-graining satisfying $M\ll D$. As discussed above, this is satisfied for all that we humans can perceive. Third, the considered projectors must evolve slowly in a suitable sense such that the unobserved microscopic degrees of freedom have time to self-average or randomize. Common examples include hydrodynamic modes, collective degrees of freedom, weakly coupled subsystems, or more generally projectors that almost commute with the total Hamiltonian (for a more technical discussion about the notion of slowness see Ref. Strasberg et al. (2023)). Fourth, the isolated system should be non-integrable.
This last point is certainly still subject to debate, not the least because a universally agreed on definition of quantum (non-)integrability is lacking. For instance, van Kampen assumed that the energy gap spectrum of the Hamiltonian $H$ is non-degenerate (apart from rare accidental degeneracies) Van Kampen (1954), the results of Ref. Gemmer and Steinigeweg (2014); Schmidtke and Gemmer (2016); Nation and Porras (2020); Strasberg et al. (2023) were based on the eigenstate thermalization hypothesis D’Alessio et al. (2016); Deutsch (2018), and here we use random matrix theory as also done in Ref. Nation and Porras (2020); Strasberg (2023). Since already classical systems tend to be chaotic (e.g., the Newtonian three-body problem), we believe that non-integrability can hardly count as an assumption in our Universe, even though it is an open question whether it is strictly necessary.
@@ III.1 Setup
Each of these blocks is in principle fully determined by the microscopic Hamiltonian (12), but we only want to capture four main features here. First, the systems $A$ and $B$ are assumed to be identical from a thermodynamic point of view. In particular, we assume that the relation between energy and temperature is the same in $A$ and $B$ such that their energy difference is proportional to their temperature difference. Second, we consider normal thermodynamic systems where the subspace corresponding to an equal energy distribution (or the same temperature according to the previous agreement) is the largest “equilibrium” subspace, i.e., $V_{0}>V_{-},V_{+}$. Third, we assume a weak interaction between $A$ and $B$, i.e., the interaction energy $\lambda H_{I}$ is supposed to be negligible (otherwise it would not be justified to restrict the discussion to a microcanonical energy window defined by the sum $E_{A}+E_{B}$ of local energies only). Fourth, we want to mimic the interaction of two generic complex (and thus non-integrable) many-body systems. According to common knowledge in statistical mechanics, this can be efficiently done by using random matrix theory Wigner (1967); Brody et al. (1981); Beenakker (1997); Guhr et al. (1998); Haake (2010); D’Alessio et al. (2016); Deutsch (2018).
@@ III.2 Results
To continue with the many worlds simulation, we now choose an equilibrium initial state with $p_{x}(0)=V_{x}/D$, for reasons that will also become clear later. Naively, one would expect nothing interesting in such a Universe as indicated in Fig. 3(b): aside from fluctuations, which are exponentially suppressed as a function of the Hilbert space dimension, nothing seems to happen. But interestingly, a completely different picture is possible: even from an equilibrium state many different interesting nonequilibrium universes can emerge. However, to warrant such a conclusion, we first of all need to ensure that the histories defined by the present coarse-graining are decoherent, i.e., they obey the DHC condition as discussed in the previous section. In the following, the evaluation of the DF is done for constant time intervals $t_{k+1}-t_{k}$ equal to the nonequilibrium relaxation time scale $\tau$ defined in Eq. (17).
@@
It is intriguing to ask whether histories or branches that look more distinct from a macroscopic point of view are characterized by stronger decoherence. Specifically, let us define the Hamming distance $d(\textbf{x},\textbf{y})$ between two histories x and y as the number of labels where they differ. For instance, for $\textbf{x}=(0,+,0,-,0)$ and $\textbf{y}=(0,-,+,0,0)$ we have $d(\textbf{x},\textbf{y})=3$. Even though, to the best of our knowledge, this question has never been asked before, it seems intuitive to expect that more distinct histories show stronger decoherence. This intuition is indeed confirmed in Fig. 6, even though the dependence on the distance $d(\textbf{x},\textbf{y})$ is rather mild: an approximate twofold increase in decoherence when going from $d(\textbf{x},\textbf{y})=1$ to $d(\textbf{x},\textbf{y})=4$. However, in more realistic scenarios one would need to take into account much longer histories. Moreover, it seems that the differences are more pronounced for larger Hilbert space dimensions $D$. Unfortunately, current numerical limitations do not allow us to draw strong conclusions from our data.
Figure 6: Average decoherence ${\langle{\epsilon}\rangle}$ as a function of the history distance $d(\textbf{x},\textbf{y})$. Here, ${\langle{\epsilon}\rangle}=\sum_{d^{\prime}=d(\textbf{x},\textbf{y})}\epsilon(\textbf{x};\textbf{y})/\#(d^{\prime})$ sums the normalized DF in Eq. (4) over all pairs of histories with a fixed distance $d^{\prime}=d(\textbf{x},\textbf{y})$, divided by the number #$\#(d^{\prime})$ of such pairs. We display results for $D=500$ (magenta circles), $D=5000$ (yellow triangles) and $D=50000$ (blue crosses) for $L=5$ and for the same nine realizations as in Figs. 4 and 5.
The overall picture that emerges from our model is depicted in Fig. 7. Figure 7 (a) shows a branching tree structure with respect to histories defined by energetic macrostates labeled by $x\in\{-,0,+\}$ for $L=3$ time steps. Importantly, we can now reason about these histories using a classical ontological model for sufficiently large $D$. To facilitate talking about it, we use Boltzmann’s entropy concept $S_{B}(x)=k_{B}\ln V_{x}$. We label transitions of the system from a lower to a higher Boltzmann entropy state (i.e., from $+$ or -- to 00) as “forward” transitions (blue dashed lines) to indicate that they comply with the conventional second law-like arrow of time when $t_{0}$ denotes the initial time. However, “backward” transitions (pink dash-dotted lines) from a higher to a lower entropy state (i.e., from 00 to $+$ or --) are also possible. Finally, transitions involving no change in Boltzmann entropy are labeled by “no arrow” (black lines). The Multiverse thus consists of different histories or branches that do not interfere and that describe universes with different entropic arrows of time (including the possibility of universes with no arrow of time and both arrows of time) determined by the question whether heat flows from hot to cold or vice versa. Note that these entropic arrows of time should be distinguished from the “branching arrow of time” (i.e., the direction in which the number of branches increases). The branching arrow is pure convention and unrelated to the perceived entropic arrow of time within the universes.
@@
For one realization of the Hamiltonian and one realization of the initial state we depict the probability for each history, which can be now interpreted classically, by the histogram below the tree-like branching structure in Fig. 7 (b). As a consistency check, we immediately see that the history $\textbf{x}=(0,0,0)$, in which the universe resides in the dominant equilibrium macrostate $x=0$, is the most probable one. Note that the probability for this history is different from the probability $(V_{0}/D)^{3}=0.6^{3}=0.216$ that one would expect by sampling the histories at time intervals comparable to the equilibration time. The time steps $\tau$ we use are clearly in the nonequilibrium regime.
An important consistency check is to ask about the probability of forward or backward arrows of time. Since the Schrödinger equation obeys time reversal symmetry and since the initial equilibrium state that we chose also does not introduce any time asymmetry, we should expect to get a time symmetric answer in this case. The answer is shown in Fig. 7 (c), (d) and (e) where we plot the probability for a forward arrow of time (blue), no arrow of time (black) and a backward arrow of time (pink) for histories of length $L=3$, $L=4$ and $L=5$, respectively. Here, histories with multiple arrows of time contribute according to their “net” arrow of time, i.e., the history $\textbf{x}=(0,+,0)$ has no overall arrow of time. Owing to the fact that we have only three macrostates, there can be no histories with two or more net forward or backward arrows of time. Thus, our model only exhibits “mini” entropic arrows of time, unable to account for our arrow of time in reality. Clearly, by including more than three energy windows in the coarse graining, we could observe longer arrows of time in this model, but their probability (when starting from an equilibrium state) would be very small albeit non-zero. This is the (in)famous Boltzmann brain paradox Albrecht and Sorbo (2004); Carroll (2020); Müller (2020). In any case, we see that the arrows are indeed symmetrically distributed in our model as it should be. Some non-visible small random fluctuations (in the third significant digit) are due to the fact that a random initial state can still have a small preference to evolve into a higher or lower entropy region, i.e., it is not perfectly symmetric under time reversal. The histograms in Fig. 7 are based on a single realization of the Hamiltonian and initial state, but we again found this behaviour to be typical.
@@
The results for the emergence of decoherence with initial energy eigenstates are shown in Figs. 8 and 9. We again observe a rather typical behaviour in Fig. 8 and a scaling law suggesting the exponential suppression of quantum effects as a function of the particle number. However, compared with Figs. 4 and 5 the variance in the datapoints is appreciably larger, i.e., there is less typicality. Moreover, the exponents are different. For $\epsilon$ we find $\alpha\approx 0.44$, which is close to the previous value of $\alpha\approx 0.5$. Instead, for $\Delta_{L}^{\text{max}}$ we find $\alpha\approx 0.2$, which is half of the value we found for Haar random initial equilibrium states. Moreover, $\Delta_{L}^{\text{max}}$ fluctuates strongly for different realizations such that the fitted scaling law (red solid line) provides only a rough orientation. Thus, the preliminary conclusions that we can draw from these data is that correlation functions of energy eigenstates behave differently from Haar random equilibrium states (which will typically overlap with many energy eigenstates), which is an interesting result in its own right. In particular, Albrecht et al. recently introduced an “eigenstate einselection hypothesis” (Appendix B in Ref. Albrecht et al. (2022)), where they claim that the “general features of the consistent histories quantities are unchanged” for energy eigenstates. On a quantitative level we see here clear evidence for the opposite, but from a qualitative point of view we observe the same: the difference between $D^{-0.4}$ and $D^{-0.2}$ might not matter in practice, say, for $D\geq 10^{100}$. Thus, the emergence of decoherence as studied here seems to be a rather robust phenomenon if we restrict the attention to slow and coarse observables of a non-integrable quantum many-body system.
@@
While the considered observable has the same coarseness, we now observe a quite different behaviour of decoherence in Figs. 10 and 11. While typicality still holds well, a much milder exponential suppression of quantum effects is observed for $\epsilon$ and histories of length $L\geq 3$, with a scaling exponent around $\alpha=0.15$ (Fig. 10). Even more drastic changes appear for $\Delta_{L}^{\text{max}}$ (Fig. 11) with an exponent $\alpha\approx 0$ for $L=5$ (indicating perhaps a power-law suppression or no suppression at all). Recalling that we have to base our conclusions on a finite (and rather small) amount of samples, a clear-cut conclusion seems not possible: exponential suppression of quantum effects, which requires $\alpha>0$, seems possible but is not warranted, and it is certainly much weaker than in the weak coupling (slow observable) regime.
@@
Notably, exponential suppression of quantum effects (with a large exponent $\alpha\approx 0.5$) still holds for the shortest length of histories with $L=2$. This seems counterintuitive, but it can be explained by recalling that we have chosen a Haar random initial equilibrium state of the form (18) with $p_{x}(0)=V_{x}/D$ in Figs. 10 and 11. The two-time DF then only probes quantum features of a two-time equilibrium correlation function. They are decoherent because, first, $\mathbb{E}[\Pi_{x_{0}}|\psi(t_{0}){\rangle\langle}\psi(t_{0})|\Pi_{y_{0}}]\sim\delta_{x_{0},y_{0}}$ where $\mathbb{E}[\dots]$ denotes a Haar random average, and second, typicality implies that most pure states $|\psi(t_{0})\rangle$ behave similarly. Indeed, for this situation an analytical proof of decoherence has been given in Ref. Strasberg et al. (2023).